# Non-relativistic transport from frame-indifferent kinetic theory

### v1

@@ I Introduction
The development of kinetic theory for rotating fluids on curved surfaces, particularly through the derivation of odd viscosity from microscopic dynamics, opens new avenues in the study of fluid behavior under rotation. Odd viscosity is a non-dissipative term that emerges in the equations governing rotating fluids, reflecting their inherent lack of time-reversal symmetry—a characteristic not accounted for in classical fluid dynamics. By applying these insights to macroscopic fluid models, such as rotating shallow water systems, researchers can incorporate odd viscosity to regularize otherwise pathological behaviors in the spectral properties of fluid waves at high wavenumbers [45, 46]. Our computation generalizes previous non-covariant computations of odd transport based on rotating reference frames [47, 48].
@@ II Basics of Newton-Cartan geometry
Newton-Cartan (NC) spacetimes can be constructed from a non-relativistic limit of pseudo-Riemannian geometries. Such a limit is usually taken either by an explicit expansion in $1/c^{2}$ [42, 43] or by a so-called lightcone reduction [54, 55, 56]. However, in this paper we will follow a more axiomatic perspective in the spirit of [38, 57].
A NC geometry is a $(d+1)$-manifold with coordinates $x^{\mu}$ with $\mu=0,1,\ldots d$, a 1-form $\tau_{\mu}$ usually called the clock form or time metric, a degenerate symmetric tensor $h^{\mu\nu}$ interpreted as an inverse spatial metric, and a covariant derivative $\nabla_{\mu}$. The degeneracy condition manifests as $\tau_{\mu}h^{\mu\nu}=0$. In addition, the derivative is required to obey $\nabla_{\mu}\tau_{\nu}=\nabla_{\mu}h^{\nu\rho}=0$. We assume that the coordinates, the time metric, and the spatial metric have units $[x^{\mu}]=L$, $[\tau_{\mu}]=T/L$, and $[h^{\mu\nu}]=1$, respectively.
Now let us introduce a set of observers with associated velocity fields $(v_{\psi})^{\mu}$, where $\psi$ labels the observer, normalized such that $\tau_{\mu}(v_{\psi})^{\mu}=1$. Any such velocity field can be expressed as
@@
with $v^{\mu}$ an arbitrary reference velocity. The reference velocity $v^{\mu}$ makes it possible to define a metric $h_{\mu\nu}$ satisfying
@@
Notice, however, that the ambiguity in the definition of $v^{\mu}$ implies that the metric $h_{\mu\nu}$ is observer-dependent. Namely, changing the reference velocity $v^{\mu}\rightarrow(v_{\psi})^{\mu}$ changes the metric as
@@
where $P^{\mu}_{\nu}=\delta^{\mu}_{\nu}-v^{\mu}\tau_{\nu}$ and $\psi^{2}=h^{\alpha\beta}\psi_{\alpha}\psi_{\beta}$. With these ingredients we find that a connection associated to the covariant derivative $\nabla_{\mu}$ has the general form [38]  [^1] We will use round brackets to denote symmetrization $A_{(\mu}B_{\nu)}=\frac{1}{2}\left(A_{\mu}B_{\nu}+A_{\nu}B_{\mu}\right)$, and square brackets for antisymmetrization $A_{[\mu}B_{\nu]}=\frac{1}{2}\left(A_{\mu}B_{\nu}-A_{\nu}B_{\mu}\right)$ respectively.
@@
with $F_{\mu\nu}$ an antisymmetric tensor with units $[F_{\mu\nu}]=T^{-1}$. A straightforward but tedious computation reveals that the observer’s acceleration $a^{\mu}\equiv v^{\nu}\nabla_{\nu}v^{\mu}$ and vorticity $\omega^{\mu\nu}\equiv 2\nabla^{[\mu}v^{\nu]}$ are given by
$a^{\mu}=-F^{\mu}\,_{\sigma}v^{\sigma}\,,~{}~{}~{}~{}\omega^{\mu\nu}=F^{\mu\nu}\,,$ (5)
where the indices are raised with the inverse spatial metric $h^{\mu\nu}$. The connection defined in Eq. (4) is in general observer-dependent. However, if we introduce a frame-dependent gauge field $m_{\mu}$ such that $F_{\mu\nu}=\partial_{\mu}m_{\nu}-\partial_{\nu}m_{\mu}$ and postulate the transformation
@@
Thus, the connection $\Gamma^{\mu}_{\alpha\beta}$ becomes frame-invariant if $d\tau=0$, which is always the case when there exists a globally define time function such that $\tau=dt$, and which is furthermore equivalent to setting the torsion $\Gamma^{\mu}_{[\alpha\beta]}$ to zero. The ambiguity expressed by Eqs. (1), (3) and (6) is known in the literature as Milne boosts. Requesting Milne boost invariance can be interpreted as the independence of physical phenomena from the observer describing them.
Finally, a volume form satisfying the Milne- and gauge-invariance requirement can be defined after introducing the non-degenerate bilinear form $G_{\mu\nu}=c^{2}\tau_{\mu}\tau_{\nu}+h_{\mu\nu}$ that allows us to define the volume form [^2] To construct $G_{\mu\nu}$ it is necessary to have a velocity scale $c$. This parameter a priori does not have to be the speed of light. However, as we will see below, observables will not depend on such a constant; therefore, we can fix it to be the speed of light.
@@
II.1 Flat Newton Cartan, Milne boosts and Galilean transformations
From the perspective of the Galilean group the presence of the gauge field is natural since mass is a conserved charge for non-relativistic matter, and therefore we can expect the presence of a gauge field that is coupled to the mass current. However, the Milne transformations are less intuitive and they do not have a counterpart for relativistic systems. As we will discuss below, Milne boosts are a necessary ingredient to obtain the Galiean group as the isometry group of the flat NC geometry.
@@ II.1 Flat Newton Cartan, Milne boosts and Galilean transformations
where £$\pounds_{\xi}$ denotes the Lie derivative along $\xi^{\mu}$. The flat NC geometry is given by
@@
and its isometry group is the set of transformations $\delta_{\chi_{K}}$ leaving the fields invariant. In particular the set of Killing generators is
@@
with commutation relations
@@
This is an extension of the Galilean algebra with a central element $M$, associated with mass. The generators correspond to time translations $(H)$, spatial translations $(P_{i})$, rotations $(R_{ij})$, Galilean boosts $(K_{i})$, and the central charge $(M)$. Such a construction allows one to have the proper field content that couples to the energy, momentum, and mass currents. Moreover, we emphasize that without the Milne boosts the algebra in Eqs. (18) would not close.
@@ III Motion of a point particle
$\dot{t}=\text{const.}\implies\tau_{\mu}\dot{x}^{\mu}(\lambda)=\text{const.}\,.$. . (20)
@@
where $m$ is the mass of the particle. Notice that under gauge transformations the Lagrangian $\mathcal{L}$ transforms as a total derivative. The affinely parametrized ($\tau_{\mu}\dot{x}^{\mu}=1$) trajectories satisfy the equations of motion
@@
where $p^{2}=h_{\mu\nu}p^{\mu}p^{\nu}$, and $p^{\mu}$ is the kinematic momentum satisfying the constraint $\tau_{\mu}p^{\mu}=m$. Notice that the right-hand side of Eq. (23) resembles a Lorentz force on the particle after interpreting $-p^{2}/2m$ as the corresponding charge. The canonical momentum reads [59, 60, 61]
@@
However, not all components of $\pi_{\mu}$ are independent [^3] This can be traced down to the reparametrization invariance of the action, which implies that the theory has a Hamiltonian constraint.. In fact, in a $(d+1)$-dimensional spacetime $\pi_{\mu}$ has only $d$ independent components due to the phase space constraint
@@
Therefore, we find it more convenient to invert Eq. (24) as
@@
and use the non-canonical pair $(x^{\mu},p^{\mu})$ as phase space coordinates with $p^{\mu}$ obeying the constraint $\tau_{\mu}p^{\mu}=m$. Thus, the time evolution of any phase space function $F(x,p)$ is computed as
@@
For generic geometries the system will not have any Noether charge; however, our ultimate goal is to construct a kinetic theory where a gas of interacting NC particles can equilibrate. In that case it is mandatory to restrict the problem to geometries with at least one time-like Killing transformation. Therefore, we assume the existence of a set of parameters $\chi_{K}=\left(\xi_{K}^{\mu}\partial_{\mu},\psi^{K}_{\mu}dx^{\mu},\Lambda_{K}\right)$ such that the NC data is invariant under the action of $\delta_{\chi_{K}}$, i.e.,
@@
In this case, we call the set of parameters $\chi_{K}$ the Killing parameters and the conditions written in Eq. (30) the Killing conditions. Since the NC data are invariant under the transformation $\chi_{K}$, then it is clear that the action in Eq. (21) is also invariant under this variation, up to a possible boundary term, and there will exist a corresponding Noether charge [62].
Let us consider an arbitrary infinitesimal coordinate transformation $x^{\mu}\rightarrow x^{\mu}+\xi^{\mu}$. The on-shell variation of a Lagrangian (see, e.g., [62]) is expressed as
$\delta_{\mathrm{on}}\mathcal{L}$ $=\xi^{\mu}\frac{\partial\mathcal{L}}{\partial x^{\mu}}+\dot{\xi}^{\mu}\frac{\partial\mathcal{L}}{\partial\dot{x}^{\mu}}=\frac{d}{d\lambda}\left(\xi^{\mu}\frac{\partial\mathcal{L}}{\partial\dot{x}^{\mu}}\right)\,.$ (31)
In the last equality, we have used the Euler-Lagrange equations. On the other hand, the symmetry variation of the Lagrangian $\mathcal{L}$ given in Eq. (21) under a Killing transformation $\chi_{K}$ reads
@@
$=-m\frac{d}{d\lambda}\Lambda_{K}\,.$ (32)
The details of this procedure are discussed in [63, 64, 65]; for a detailed calculation, see Appendix B.1. Now, if we replace $\xi^{\mu}$ by $\xi_{K}^{\mu}$ in the on-shell variation in Eq. (31), then the left-hand sides of both variations are equal [62]. Comparing Eqs. (31) and (32) we get the conserved charge as
$Q_{K}=\xi_{K}^{\mu}\pi_{\mu}+m\Lambda_{K}\,.$ (33)
Note that $Q_{K}$ is a scalar that is both Milne-invariant and gauge-invariant.
In addition, we find it convenient to fix units such that Planck’s, Boltzmann’s, and the gravitational constants take values $h=k_{B}=G_{N}=1$. In these units, the relevant physical quantities of the problem can be expressed in powers of length only and read $[mass]=L^{-1/3}$, $[momentum]=L^{-1}$, and $[energy]=[temperature]=[T^{-1}]=L^{-5/3}$.

### v2

@@ I Introduction
The development of kinetic theory for rotating fluids on curved surfaces, particularly through the derivation of odd viscosity from microscopic dynamics, opens new avenues in the study of fluid behavior under rotation. Odd viscosity is a non-dissipative term that emerges in the equations governing rotating fluids, reflecting their inherent lack of time-reversal symmetry—a characteristic not accounted for in classical fluid dynamics. By applying these insights to macroscopic fluid models, such as rotating shallow water systems, researchers can incorporate odd viscosity to regularize otherwise pathological behaviors in the spectral properties of fluid waves at high wavenumbers [45, 46, 47]. Our computation generalizes previous non-covariant computations of odd transport based on rotating reference frames [48, 49].
@@ II Basics of Newton-Cartan geometry
Newton-Cartan (NC) spacetimes can be constructed from a non-relativistic limit of pseudo-Riemannian geometries. Usually, this limit is taken either by an explicit expansion in $1/c^{2}$ [42, 43] or by a so-called light-cone reduction [55, 56, 57]. However, in this paper, we will follow a more axiomatic perspective in the spirit of [38, 58].
A NC spacetime is a $(d+1)$-manifold with coordinates $x^{\mu}$ with $\mu=0,1,\ldots d$ and:
  • a 1-form $\tau_{\mu}$ usually called the clock form or time metric, such that $\tau_{\mu}(x_{2}^{\mu}-x_{1}^{\mu})$ measures the passage of time between two infinitesimally close points $x_{1}^{\mu}$ and $x_{2}^{\nu}$;
  • a degenerate symmetric tensor $h^{\mu\nu}$, with signature $(0,1,\ldots,1)$, that satisfies $\tau_{\mu}h^{\mu\nu}=0$, interpreted as inverse spatial metric;
  • a covariant derivative $\nabla_{\mu}$ that acts on a tensor $T_{\alpha}^{~\beta}$ as
    $\nabla_{\mu}T_{\alpha}^{~\beta}=\partial_{\mu}T_{\alpha}^{~\beta}+\Gamma^{\beta}_{\lambda\mu}T_{\alpha}^{~\lambda}-\Gamma^{\lambda}_{\alpha\mu}T_{\lambda}^{~\beta}.$ (1)
    As in general relativity, we require it to be compatible with the geometric data, $\nabla_{\mu}\tau_{\nu}=\nabla_{\mu}h^{\nu\rho}=0$.
Throughout the paper, we raise indices using the inverse spatial metric, eg. $T^{\alpha\beta}=h^{\alpha\gamma}T_{\gamma}^{~\beta}$. Note that since $h^{\mu\nu}$ is degenerate, raising an index is not invertible; therefore, the distinction between upper-index and lower-index components of tensors is more significant than in relativistic physics. We assume that the coordinates, the time metric, and the spatial metric have units $x^{\mu}]=\mathrm{length}$, $\tau_{\mu}]=\mathrm{time}/\mathrm{length}$, and $h^{\mu\nu}]=1$, respectively.
Now let us introduce a family of observers labelled by $\psi$ with an associated velocity field $(v_{\psi})^{\mu}$, normalized such that $\tau_{\mu}(v_{\psi})^{\mu}=1$. Any such velocity field can be expressed as a Milne boost of a chosen reference velocity
@@ A NC spacetime is a $(d+1)$-manifold with coordinates $x^{\mu}$ with $\mu=0,1,\ldots d$ and:
with $v^{\mu}$ an arbitrary reference velocity satisfying $\tau_{\mu}v^{\mu}=1$. The reference velocity $v^{\mu}$ makes it possible to define a metric $h_{\mu\nu}$ satisfying
@@
Milne boosts express a gauge freedom in the choice of geometric data in NC geometry. Specifically, they reflect the freedom to redefine the reference velocity without altering the underlying physical structure. This gauge freedom is implemented via transformations given by Eq. (2), and leads to a corresponding transformation of the spatial metric. As a result, the spatial metric $h_{\mu\nu}$ becomes observer-dependent, transforming under a Milne boost as:
@@
where $P^{\mu}_{\nu}=\delta^{\mu}_{\nu}-v^{\mu}\tau_{\nu}$ is the spatial projector and $\psi^{2}=h^{\alpha\beta}\psi_{\alpha}\psi_{\beta}$. This structure captures the fact that Newton–Cartan geometry is not fixed by a unique choice of $v^{\mu}$, but instead allows for an entire gauge orbit of related geometric data connected by Milne boosts. Requesting Milne boost invariance of the theory can be interpreted as the independence of physical phenomena from the observer describing them.
As shown in [38], the affine connection $\Gamma^{\mu}_{\alpha\beta}$ must transform accordingly to preserve Milne invariance. With these ingredients we find that the connection $\Gamma^{\mu}_{\alpha\beta}$ has the general form [38] [^1] We will use round brackets to denote symmetrization $A_{(\mu}B_{\nu)}=\frac{1}{2}\left(A_{\mu}B_{\nu}+A_{\nu}B_{\mu}\right)$, and square brackets for antisymmetrization $A_{[\mu}B_{\nu]}=\frac{1}{2}\left(A_{\mu}B_{\nu}-A_{\nu}B_{\mu}\right)$ respectively.
@@
with $F_{\mu\nu}$ an antisymmetric tensor with units $F_{\mu\nu}]=\mathrm{time}^{-1}$. A straightforward but tedious computation reveals that the geodesic acceleration $a_{v}^{\mu}\equiv v^{\nu}\nabla_{\nu}v^{\mu}$ and vorticity $\omega_{v}^{\mu\nu}\equiv 2\nabla^{[\mu}v^{\nu]}$ are given by
$a_{v}^{\mu}=-F^{\mu}\,_{\sigma}v^{\sigma}\,,~~~~\omega_{v}^{\mu\nu}=F^{\mu\nu}\,$, (6)
where the indices are raised with $h^{\mu\nu}$. The connection defined in Eq. (5) is in general observer-dependent, but if we introduce a frame-dependent gauge field $m_{\mu}$ such that $F_{\mu\nu}=\partial_{\mu}m_{\nu}-\partial_{\nu}m_{\mu}$ and postulate the transformation
@@
Thus, the connection $\Gamma^{\mu}_{\alpha\beta}$ becomes observer-invariant if $d\tau=0$, which is always the case when there exists a globally defined time function such that $\tau=dt$, and which is furthermore equivalent to setting the torsion $\Gamma^{\mu}_{[\alpha\beta]}$ to zero.
Finally, a volume form satisfying the Milne- and gauge-invariance requirement can be defined after introducing the non-degenerate tensor $G_{\mu\nu}=c^{2}\tau_{\mu}\tau_{\nu}+h_{\mu\nu}$ that allows us to define the volume form [^2] To construct $G_{\mu\nu}$ it is necessary to have a velocity scale $c$. This parameter a priori does not have to be the speed of light. However, as we will see below, observables will not depend on such a constant; therefore, we can fix it to be the speed of light.
@@
II.1 Flat NC geometry, Milne boosts and Galilean transformations
From the perspective of the Galilean group, the presence of the gauge field $m_{\mu}$ is natural since mass is a conserved charge for non-relativistic matter, and therefore we can expect the presence of a gauge field that couples to the mass current. However, the Milne transformations are less intuitive and they do not have a counterpart for relativistic systems. As we will discuss below, Milne boosts are a necessary ingredient to obtain the Galilean group as the isometry group of the flat NC geometry.
@@ II.1 Flat NC geometry, Milne boosts and Galilean transformations
where $\pounds_{\xi}$ denotes the Lie derivative along $\xi^{\mu}$. These transformations form an algebra as $\left[\delta_{\chi_{1}},\delta_{\chi_{2}}\right]=\delta_{\chi_{[12]}}$, where $\chi_{1}=\left(\xi^{\mu}_{1},\psi_{\mu}^{1},\Lambda_{1}\right)$ and $\chi_{2}=\left(\xi^{\mu}_{2},\psi_{\mu}^{2},\Lambda_{2}\right)$ are two sets of symmetric parameters, and $\chi_{[12]}=\left(\xi^{\mu}_{[12]},\psi_{\mu}^{[12]},\Lambda_{[12]}\right)$ with
$\xi^{\mu}_{[12]}$ $=\pounds_{\xi_{1}}\xi_{2}\,$, (15)
$\psi_{\mu}^{[12]}$ $=\pounds_{\xi_{1}}\psi_{\mu}^{2}-\pounds_{\xi_{2}}\psi_{\mu}^{1}\,$, (16)
$\Lambda_{[12]}$ $=\pounds_{\xi_{1}}\Lambda_{2}-\pounds_{\xi_{2}}\Lambda_{1}\,.$ (17)
In particular, the flat NC geometry is given by
@@
and its isometry group is the set of transformations $\delta_{\chi}$ leaving the fields invariant. In particular, the set of Killing generators is
@@
satisfying the commutation relations of the so-called Bargmann algebra [^3] This is an extension of the Galilean algebra with a central element $M$, associated with mass.
@@
The generators correspond to time translations $(H)$, spatial translations $(P_{i})$, rotations $(R_{ij})$, Galilean boosts $(K_{i})$, and the central charge $(M)$. We note that the index on the generators serves as a label and does not correspond to a spatial component. Such a construction allows one to have the proper field content that couples to the energy, momentum, and mass currents. Moreover, we emphasize that without the Milne boosts the algebra in Eqs. (22) would not close.
II.2 Thermal NC spacetime
In [58] NC spacetimes were generalized by adding to the manifold an extra timelike vector field $u^{\mu}$, normalized as $u^{\mu}\tau_{\mu}=1$. As we will see later, at finite density such a field can always be defined, in which case $u^{\mu}$ describes the velocity of the gas. For such systems we could argue that the “Milne symmetry” is spontaneously broken since $u^{\mu}$ introduces a distinguished Milne frame with velocity $v_{u}^{\mu}=u^{\mu}$. Notice that this frame is connected to an arbitrary one via a Milne boost $v^{\mu}-v_{u}^{\mu}=h^{\mu\nu}\psi_{\nu}$. The presence of $u^{\mu}$ allows us to define the spatial metric $g_{\mu\nu}$ and $U(1)$ gauge field $A_{\mu}$
$A_{\mu}$ $=m_{\mu}+u_{\mu}-\frac{u^{2}}{2}\tau_{\mu}\,$, (24)
where $u_{\mu}=h_{\mu\nu}u^{\nu}$ and $u^{2}=h_{\mu\nu}u^{\mu}u^{\nu}$. The objects $g_{\mu\nu}$ and $A_{\mu}$ are precisely equal to $h_{\mu\nu}$ and $m_{\mu}$ evaluated in the distinguished frame $v_{u}^{\mu}$. A straightforward calculation shows that even though the definitions (23) and (24) involve $h_{\mu\nu}$ and $m_{\mu}$, $g_{\mu\nu}$ and $A_{\mu}$ do not transform under Milne boosts as described by Eqs. (4), (7) and are therefore defined unambiguously. We furthermore introduce the distinguished spatial projector and an affine connection corresponding to $v_{u}^{\mu}$ as, respectively, [58]
with
f=dA.f=dA\,. (27)
A comparison with Eq. (8) reveals that $\bar{\Gamma}^{\mu}_{\alpha\beta}=\Gamma^{\mu}_{\alpha\beta}$ in the torsionless case. Using the connection $\bar{\Gamma}^{\mu}_{\alpha\beta}$, which is both gauge- and Milne-invariant, the covariant derivative of the velocity field can be decomposed as
$\bar{\nabla}_{\mu}u^{\nu}=\tau_{\mu}a^{\nu}+\frac{1}{2}g_{\mu\alpha}\omega^{\alpha\nu}+g_{\mu\alpha}\sigma^{\alpha\nu}+\frac{1}{d}\bar{P}_{\mu}^{\nu}\theta\,$, (28)
$a^{\mu}$ $\equiv u^{\nu}\bar{\nabla}_{\nu}u^{\mu}=-f^{\mu}\,_{\nu}u^{\nu}\,$, (29)
$\sigma^{\mu\nu}$ $\equiv\bar{\nabla}^{\langle\mu}u^{\nu\rangle}\,$, (31)
$\theta$ $\equiv\bar{\nabla}_{\mu}u^{\mu}\,$, (32)
The new structure added to the NC manifold may seem unnecessary. However, in what follows we will prove that the existence of the notion of thermal equilibrium is only possible on this class of NC spaces. Therefore, we will call them Thermal Newton-Cartan (TNC) manifolds.
@@ III Motion of a point particle
$\dot{t}=\text{const.}\implies\tau_{\mu}\dot{x}^{\mu}(\lambda)=\text{const.}\,.$ (35)
@@
where $m$ is the mass of the particle. Notice that under gauge transformations, the Lagrangian $\mathcal{L}$ transforms as a total derivative. The affinely parametrized ($\tau_{\mu}\dot{x}^{\mu}=1$) trajectories satisfy the (non-geodesic) equations of motion
@@
where $p^{2}=h_{\mu\nu}p^{\mu}p^{\nu}$, and $p^{\mu}$ is the kinematic momentum satisfying the constraint $\tau_{\mu}p^{\mu}=m$. Notice that the right-hand side of Eq. (38) resembles a Lorentz force on the particle after interpreting $p^{2}/2m$ as the corresponding charge. The canonical momentum reads [60, 61, 62]
@@
However, not all components of $\pi_{\mu}$ are independent [^4] This can be traced to the reparametrization invariance of the action, which implies that the theory has a Hamiltonian constraint.. In fact, in a $(d+1)$-dimensional spacetime, $\pi_{\mu}$ has only $d$ independent components due to the phase space constraint
@@
Therefore, we find it more convenient to invert Eq. (39) (making use of $\tau_{\mu}p^{\mu}=m$) as
@@
and use the non-canonical pair $(x^{\mu},p^{\mu})$ as phase space coordinates, with $p^{\mu}$ obeying the constraint $\tau_{\mu}p^{\mu}=m$.
Thus, the time evolution of any phase space function $F(x,p)$ is computed as
@@
For generic geometries, the system will not have any Noether charge; however, our ultimate goal is to construct a kinetic theory, where a gas of interacting NC particles can equilibrate. In that case, it is mandatory to restrict the problem to geometries with at least one time-like Killing transformation. Therefore, we assume the existence of a set of parameters $\chi_{K}=\left(\xi_{K}^{\mu}\partial_{\mu},\psi^{K}_{\mu}dx^{\mu},\Lambda_{K}\right)$ such that the NC data is invariant under the action of $\delta_{\chi_{K}}$, i.e.,
@@
Since the NC data are invariant under the transformation $\chi_{K}$, then it is clear that the action in Eq. (36) is also invariant under this variation, up to a possible boundary term, and there will exist a corresponding Noether charge [63].
Let us consider an arbitrary path deformation $x^{\mu}(\lambda)\rightarrow x^{\mu}(\lambda)+\tilde{\xi}^{\mu}(x(\lambda))$. The on-shell variation of the Lagrangian (see, e.g., [63]) is expressed as
$\delta_{\mathrm{on}}\mathcal{L}$ $=\tilde{\xi}^{\mu}\frac{\partial\mathcal{L}}{\partial x^{\mu}}+\dot{\tilde{\xi}}^{\mu}\frac{\partial\mathcal{L}}{\partial\dot{x}^{\mu}}=\frac{d}{d\lambda}\left(\tilde{\xi}^{\mu}\frac{\partial\mathcal{L}}{\partial\dot{x}^{\mu}}\right)\,$, (46)
In the last equality, we have used the Euler–Lagrange equations. On the other hand, the symmetry variation of the Lagrangian $\mathcal{L}$ given in Eq. (36) under a Killing transformation $\xi_{K}^{\mu}$ reads
@@
$=-m\frac{d}{d\lambda}\Lambda_{K}\,$, (47)
The details of this procedure are discussed in [64, 65, 66, 67]; for a detailed calculation, see Appendix B.1. Now, if we replace $\tilde{\xi}^{\mu}$ by $\xi_{K}^{\mu}$ in the on-shell variation in Eq. (46), then the left-hand sides of both variations are equal [63]. Comparing Eqs. (46) and (47), we obtain the conserved charge as
$Q_{K}=\xi_{K}^{\mu}\pi_{\mu}+m\Lambda_{K}\,$, (48)
Note that $Q_{K}$ is a scalar that is both Milne- and gauge-invariant.
In addition, we find it convenient to fix units such that Planck’s, Boltzmann’s, and the gravitational constants take values $h=k_{B}=G_{N}=1$. In these units, the relevant physical quantities of the problem can be expressed in powers of length only and read $mass]=\mathrm{length}^{-1/3}$, $momentum]=\mathrm{length}^{-1}$, and $energy]=[temperature]=[time]^{-1}=\mathrm{length}^{-5/3}$.